OpenStudy (anonymous):

How to solve if x^2 congruent to 7(mod11) has solutions? I am sure there is something painfully obvious that I'm not seeing..

OpenStudy (primeralph):

Yes, the tip of your nose.

OpenStudy (blurbendy):

if a ≡ b(mod m) then by the def. of congruent modulo m | (a - b) then, if you let x = even integers x = 2k then, if you let x = odd integers x = 2k + 1 then play around with these definitions, and the stars should align (hopefully)

OpenStudy (blurbendy):

a ≡ b(mod m) is also equivalent to saying a % m = b % m but that should be obvious, and i dont think that definition will be as helpful.

OpenStudy (primeralph):

@sarahusher Got it ?

OpenStudy (anonymous):

i think maybe legendre is the way to go

OpenStudy (primeralph):

@satellite73 Why not just standard definitions?

OpenStudy (anonymous):

if \(\left(\frac{7}{11}\right)=-1\) then \(7\) is a quadratic residue if i recall correctly

OpenStudy (anonymous):

No :( I'm still not getting it in the legendre ^^ if it =-1 then it is a quadratic non residue

OpenStudy (anonymous):

not sure how you go about finding the existence of square roots modulo a prime by standard definitions

OpenStudy (anonymous):

oh right sorry you want 1 not -1

OpenStudy (primeralph):

@satellite73 Well, I do; but lets have it your way.

OpenStudy (anonymous):

kk i will listen

OpenStudy (anonymous):

Oh okay, I get it in terms of the Legendre symbol now, thanks @satellite73 But I would also like to understand in the terms discussed above too, if that's okay? I want to make sure I fully understand it..

OpenStudy (anonymous):

btw my calculation gives \(-1\)

OpenStudy (anonymous):

that is \(7^5\equiv -1(\text{ mod }11)\)

OpenStudy (anonymous):

and hello myininaya!!

OpenStudy (blurbendy):

Legendre, awesome. learn something new everyday.

OpenStudy (anonymous):

Okay, yep I get -1 as well Could someone explain to me the above method by using odd/even integers?

OpenStudy (blurbendy):

if we said 11 | (a - b) going off the defintiion 11 | (2k)^2 - 7 where we are letting x be even integers 11 | 4k^2 - 7 11 | 4(k^2 - 7/4) 11 | 4(k^2 - 7/4) so the question becomes, does 11 divide (a-b) when a -b is a multiple of 4? we find that if k = \[\frac{ 3 }{ \sqrt{2} } , \frac{ -3 }{ \sqrt{2} }\]

OpenStudy (blurbendy):

then, 11 | (a - b) is true, so perhaps that is sufficient to say x^2 congruent to 7 mod 11 has solutions

OpenStudy (blurbendy):

A similar approach would then be done with odds

OpenStudy (primeralph):

@blurbendy No argument, @satellite73 gets the job done. You haven't seen my work yet, so you can't really compare can you?

OpenStudy (blurbendy):

touche monsieur!